OpenStudy (anonymous):

A large plywood box has a volume of 1008 ft3. Its length is 16 ft greater than its height, and its width is 9 ft less than its height. What are the dimensions of the box?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

first, define all variables let L be length W be width H be height since the problem pertains to the volume of a box the formula for the volume of a box is V= L W H at this point, determine the equations of the relations in the word problem length is 16 ft greater than its height width is 9 ft less than its height once you has these substitute the values into the volume equation, solve for height, then figure out the length and width

OpenStudy (anonymous):

need more help

OpenStudy (anonymous):

at this point, determine the equations of the relations in the word problem length is 16 ft greater than its height width is 9 ft less than its height what are the equations

OpenStudy (anonymous):

using the variables defined above

OpenStudy (anonymous):

i found -- 16, 7 n 9 but m not sure which is wht

OpenStudy (anonymous):

yes when multiplied together it equals 1008, but those are not the actual values

OpenStudy (anonymous):

length is 16 ft greater than its height L = 16 + H L= 16 + H

OpenStudy (anonymous):

write this as an equation width is 9 ft less than its height